40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1323     Entry time: Thu Feb 19 04:16:17 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: Locking status 
Rob, Yoichi

We checked the CM-MC cross over just before turning off the moving zero.
There was a slight bump in the gain of the MC_L loop at (I believe) the optical spring freq. (~400Hz) just below 0 dB. The phase margin there was very thin.
Removing the moving zero will increase the bump more and make the loop unstable.
Rob suggested to increase the AO gain a bit more.

To see if the AO path is really working, I connected the OUT2 of the MC board to a spare DAQ channel (C1:PEM-OSA_APTEMP).
I confirmed that the PO_DC signal is actually coming to the AO path input of the MC board.
I also hooked up the SR785 to the A excitation channel of the common mode board, so that we can measure the loop gain of the AO path.
After these preparation, the lock acquisition process became somewhat unstable. The ifo loses lock randomly at various places in the lock acquisition steps.
So, as of 4:00 am, I have not gotten a chance to try Rob's suggestion nor the TF measurement with SR785 yet.
I will continue the work tomorrow (i.e. tonight ??).

ELOG V3.1.3-